# A Midsummer Meme’s Dream: Investigating Market Manipulations in the Meme Coin Ecosystem

### v1

@@ Abstract.
From viral jokes to a billion-dollar phenomenon, meme coins have become one of the most popular segments in cryptocurrency markets. Unlike utility-focused crypto assets like Bitcoin or Ethereum, meme coins derive value primarily from community sentiment, making them vulnerable to manipulation. This study presents a cross-chain analysis of the meme coin ecosystem, examining 34,988 tokens across Ethereum, BNB Smart Chain, Solana, and Base. We characterize the tokenomics of meme coins and track their growth in a three-month longitudinal analysis. We discover that among high-return tokens (¿100%), an alarming 82.6% show evidence of extensive use of artificial growth strategies designed to create a misleading appearance of market interest. These include wash trading and a form of manipulation we define as Liquidity Pool-Based Price Inflation (LPI), where small strategic purchases trigger dramatic price increases. We also find evidence of schemes designed to profit at the expense of investors, such as pump and dumps and rug pulls. In particular, most of the tokens involved had previously experienced wash trading or LPI, indicating how initial manipulations often set the stage for later exploitation. These findings reveal that manipulations are widespread among high-performing meme coins and suggest that their dramatic gains are often likely driven by coordinated efforts rather than natural market dynamics.
Blockchain, Meme Coins, Market Manipulations
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacyccs: Security and privacy Economics of security and privacyccs: General and reference Measurementccs: Information systems Information retrieval
1. Introduction
@@
The explosive rise of meme coins represents one of the most impressive trends in cryptocurrency markets. What started as a joke with Dogecoin in 2013 (Nani, 2022) has evolved into a serious market worth over \$117 billion by 2024 (CoinMarketCap, 2025c). Top meme coins like Dogecoin (\$53 billion), Shiba Inu (\$15 billion), and Pepe (\$8 billion) now rank among the most valuable cryptocurrencies, rivaling tokens created for functional use. Their appeal lies in the potential for extreme returns: Pepe has yielded gains of up to 400,000%, while Shiba Inu reached an incredible 80,000,000% at its peak (Coin, 2024). These extraordinary profits have attracted waves of investors hoping to find the next meme coin before its price explodes. This surge in interest has led to an unprecedented proliferation of meme coins, largely driven by platforms like pump.fun, which allow users to create tokens with minimal technical skills (Boaz Sobrado, 2025).
Despite their popularity, meme coins present unique vulnerabilities that distinguish them from other cryptocurrencies. Unlike utility tokens, meme coins rely exclusively on perceived interest to maintain their value (Nani, 2022), creating strong incentives for manipulations. Token creators and early investors can use artificial growth strategies to generate illusory activity. After attracting enough investors, these manipulators can execute profit extraction schemes, like pump and dumps (Xu and Livshits, 2019) or rug pulls (Zhou et al., 2024), to capitalize at the expense of investors. These manipulations make meme coin investments particularly risky, as trading activities may be fabricated to lure investors into projects designed to extract value from them.
@@
This paper presents a comprehensive cross-chain analysis of the meme coin ecosystem, examining the economic characteristics of these tokens and the prevalence of market manipulations within high-performing coins. We collect data on 34,988 tokens from established cryptocurrency aggregators, simulating the information environment encountered by typical investors, and focus our research on four major blockchains: Ethereum, BNB Smart Chain (BSC), Solana, and Base. Our investigation begins with a detailed characterization of meme coin tokenomics across different blockchains, examining their price structures, supply strategies, and market capitalization patterns. We then perform a three-month longitudinal analysis of price growth, identifying 707 high-return tokens (¿100%). We uncover an alarming 82.6% of these high-performing meme coins exhibit clear evidence of market manipulation, including artificial growth strategies such as recurrent wash trading and what we define as Liquidity Pool-Based Price Inflation (LPI), an operation where strategic purchases with minimal capital investment trigger dramatic price increases. Furthermore, we find profit extraction schemes designed to capitalize on artificially generated interest, with pump and dumps and rug pulls executed after calculated waiting periods. Most significantly, we discover that 62.9% of tokens subjected to these exploitative operations had previously undergone wash trading or LPI operations, revealing how initial manipulations often serve as preliminary stages in broader exploitation strategies. We also find this pattern among meme coins delisted by established aggregators: tokens first use artificial growth tactics to meet listing requirements, then leverage their newfound visibility to execute exit strategies, maximizing profits before disappearing from the ecosystem. This connection between manipulation tactics is even more pronounced in delisted tokens, where 86.67% of those subjected to profit extraction had previously undergone artificial growth operations. These findings highlight the prevalence of manipulation strategies across high-return meme coins and demonstrate the need for improved investor protection mechanisms.
@@
  • Cross-chain Characterization: To the best of our knowledge, we provide the first comprehensive cross-chain analysis of meme coins, examining 34,988 tokens across Ethereum, BSC, Solana, and Base. We find blockchain-specific patterns in temporal evolution, price structures, supply strategies, and market capitalization distributions and perform a three-month longitudinal study of meme coin price growth.
@@
  • Artificial Growth Detection: We detect strategies employed to create misleading appearances of market interest, including a form of manipulation we define as Liquidity Pool-Based Price Inflation (LPI). We find 82.6% of high-performing tokens exhibit concrete indicators of market manipulation, such as concentrated ownership structures, wash trading, or LPI activities, suggesting that the impressive gains seen in meme coins are often due to coordinated manipulation rather than natural market behavior.
  • Profit Extraction Manipulations Analysis: We identify techniques used to extract profits from tokens, including pump and dumps and rug pulls. We discover how these exploitation tactics in 62.9% of high-return involved tokens follow earlier growth manipulations (rising to 86.67% in delisted tokens), indicating how initial manipulations often set the stage for later exploitation.
2. Related Work
Memes and Meme Coins. Internet memes, defined by Shifman (Shifman, 2013) as “units of popular culture that are circulated, imitated, and transformed by individual Internet users,” have become significant cultural phenomena. Researchers studied meme propagation patterns (Zannettou et al., 2018), tracked their evolution using deep learning (Beskow et al., 2020), and analyzed how harmful content emerges through visual-textual combinations (Qu et al., 2023).
@@ 2. Related Work
The intersection of meme culture and cryptocurrencies created meme coins, which differ from traditional cryptocurrencies through their community-driven nature, humor incorporation, and meme culture origins (Stencel, 2023; Nani, 2022). In particular, the U.S. Securities and Exchange Commission (SEC) defines meme coins as a “type of crypto asset inspired by internet memes, characters, current events, or trends for which the promoter seeks to attract an enthusiastic online community to purchase the meme coin and engage in its trading“ (Securities and Commission, 2025), a definition we adopt in this study. Kim et al. (Kim and Park, 2024) revealed how social media drive meme coin valuation, while Belcastro et al. (Belcastro et al., 2023) demonstrated that meme coin price movements require specialized prediction approaches. Long et al. introduced CoinCLIP (Long et al., 2024a), a framework leveraging visual-language models to classify meme coin viability, and released the Coin-Meme dataset (Long et al., 2024b) containing multimodal data from Solana-based meme coins.
@@
Market Manipulations on Blockchains. The cryptocurrency ecosystem, characterized by its relative lack of regulatory oversight, has become susceptible to various forms of market manipulation (Gandal et al., 2018; Krafft et al., 2018; Cernera et al., 2023b; Daian et al., 2020; Cernera et al., 2024). These manipulative practices undermine market integrity and pose significant risks to investors.
@@
Wash Trading is a manipulation that aims to create an illusion of market activity by generating artificial volume. Victor and Weintraud (Victor and Weintraud, 2021) found that up to 35% of ERC-20 token volume on certain exchanges consisted of wash trades. Similarly, Le Pennec et al. (Le Pennec et al., 2021) proposed machine learning approaches to identify wash trading activities across different blockchain networks. This manipulation technique has extended to non-fungible token (NFT) markets as well, with von Wachter et al. (von Wachter et al., 2022) and La Morgia et al. (La Morgia et al., 2023a) documenting significant wash trading activity in NFT marketplaces.
@@
Pump and dump schemes represent one of the most prevalent manipulation strategies in cryptocurrency markets. These schemes involve inflating an asset’s price through misleading statements and coordinated buying (the ”pump”), followed by selling the overvalued assets (the ”dump”) once enough investors have been lured in (Kamps and Kleinberg, 2018). Xu and Livshits (Xu and Livshits, 2019) developed detection mechanisms for pump and dump operations on cryptocurrency exchanges, identifying over 300 instances of such manipulations. Complementing this work, other studies (Hamrick et al., 2019; La Morgia et al., 2023c, 2020) found that pump announcements typically generated average returns of 65% for organizers, mostly at the expense of unsuspecting investors. These works also examined the communication platforms used to coordinate such schemes, identifying Telegram and Discord as the main hubs, a finding also supported by Telegram-focused research (La Morgia et al., 2023b, [n. d.], 2025).
@@
Lastly, rug pulls have become a harmful exit scam in decentralized finance (DeFi), where developers abandon projects after attracting investors. Zhou et al. (Zhou et al., 2024) analyzed 201 rug pulls and developed tools to detect associated risks, while Cernera et al. (Cernera et al., 2023b) identified ”1-day rug pulls” that generated \$240 million in profits, primarily on BNB Smart Chain.
This study investigates whether the extreme profitability observed in some high-performing meme coins arises from organic market dynamics or is driven by coordinated manipulations.
3. Methodology
@@
To capture a comprehensive view of the meme coin ecosystem, we design a methodology consisting of three steps. We first identify tokens labeled as meme coins by established cryptocurrency platforms, then expand our dataset by collecting newly created tokens from additional sources, and finally apply classification techniques to distinguish meme coins from other cryptocurrency assets.
3.1. Data Collection
@@
The meme coin ecosystem spans multiple blockchains, including well-known tokens and emerging projects. To capture this diversity, we adopt a blockchain-agnostic data collection strategy, sourcing information from mainstream crypto tracking platforms and specialized decentralized exchange (DEX) aggregators. This cross-chain approach enables us to observe how meme coin creators leverage different blockchains, track migration patterns, and identify platform preferences, uncovering cross-chain trends and insights that would be missed in a single-chain analysis.
3.1.1. Known Meme Coins
@@
We start with CoinMarketCap (CoinMarketCap, 2025d) and CoinGecko (CoinGecko, 2025b), two leading cryptocurrency tracking platforms that maintain specialized meme token categories. CoinGecko defines meme coins as “Tokens with no intrinsic value, but with a strong community that fosters social interaction” (Coingecko, 2024), while CoinMarketCap describes them as “a type of cryptocurrency that originated from internet memes. They are often created as a joke or for the purpose of satire… Some of these coins have gained popularity due to their community-driven nature and the humorous themes they embody” (CoinMarketCap, 2025e). Both definitions align with the one provided by the SEC (Securities and Commission, 2025), which we adopt as the formal definition for our work. Using their free APIs (CoinMarketCap, 2025a; CoinGecko, 2025a), we collect data on 2,930 and 2,107 meme coins, respectively. Additionally, we utilize Gecko Terminal (CoinGecko, 2025c), which categorizes tokens by meme trends. This approach reflects the definition of meme coins as assets inspired by Internet trends and culture (Securities and Commission, 2025). We develop a custom scraper to collect data on 4,735 meme coins from pools tagged with relevant meme categories, including animal, dog, cat, inu, and pump.fun (a platform for meme coin creation).
By aggregating the results of these three services and removing duplicates, we obtain a total of 8,852 distinct meme coins.
3.1.2. Searching other Meme Coins
@@
To expand our dataset, we implement a second approach focusing on newly created tokens. We develop scrapers for two DEX aggregators: DexScreener (Dexscreener, 2025) and CoinSniper (CoinSniper, 2025), which track real-time data on emerging tokens across decentralized exchanges. These platforms differ from CoinMarketCap and CoinGecko in their listing criteria. While the latter require stricter conditions, such as functional websites and listing on integrated exchanges (CoinGecko, 2025d; CoinMarketCap, 2025b), DEX aggregators implement more permissive policies: DexScreener automatically indexes tokens immediately upon the creation of the liquidity pool and the initial transaction activity (Screener, 2024), while CoinSniper requires only a basic application process, after which the projects appear immediately on their New Listings Page without further verification (CoinSniper.Net, 2025). Our scrapers collected 30,771 tokens from DexScreener and 44,611 from CoinSniper, resulting in 65,021 unique tokens.
@@
Although these platforms are commonly used to monitor meme coins due to their focus on newly launched tokens on DEXs, they could list a wide range of token types, not just meme coins. This makes it necessary to define criteria for distinguishing meme coins from conventional cryptocurrencies, utility tokens, and other digital assets. To this end, we strictly adhere to the formal definition established by the SEC, which characterizes meme coins as ”crypto assets inspired by internet memes, characters, current events, or trends” (Securities and Commission, 2025). Our analysis revealed that these cultural and thematic associations are consistently embedded within token names, creating a distinctive linguistic pattern that can be systematically identified. Using this insight, we examine the names of known meme coins collected in Sec. 3.1.1 to identify common keywords. These keywords then served as markers for detecting additional meme coins within our larger dataset.
Name-Based Classification
@@
To identify meme-related keywords, we apply Term Frequency–Inverse Document Frequency (TF-IDF) analysis to our verified meme token dataset. This technique highlights terms by assigning higher weights to words that appear frequently in meme coin names but are uncommon in general cryptocurrency naming (Ramos et al., 2003). First, in the preprocessing phase, we convert all token names to lowercase and remove special characters. Next, we tokenize these names into individual words (unigrams) and calculate TF-IDF scores for each word. To find the ideal number of keywords for our analysis, we use the elbow method, which identifies where additional keywords offer limited value in terms of relevance (Bholowalia and Kumar, 2014). As shown in Fig. 7 in the Appendix, the elbow analysis suggests 150 keywords as an optimal cutoff point. We then manually review these keywords, removing general cryptocurrency terms (like BSC, ETH, SOL) that appear because of cross-chain versions of the same token (fe.g., Dogecoin (SOL), Dogecoin (BSC)). This refinement process results in 126 meme-specific keywords. In particular, the final keyword set reveals several prominent patterns. Terms associated with successful meme coins show high frequency: dog appear in 7.24% of names, reflecting the influence of market leaders Dogecoin and Shiba-inu; cat in 6.92% of names; inu in 4.82%; pepe in 3.82%; and doge in 3.30%. We also find keywords related to contemporary trends, such as AI (4.96%) and Trump (2.39%).
We then apply this keyword-based classification to the dataset from Sec. 3.1.2, using the same preprocessing steps. We identify tokens as meme coins if they contain one or more of the 126 meme-specific keywords, resulting in the detection of 26,557 meme coins.
Pump.fun-Based Classification
Pump.fun, a Solana-based platform specialized in meme coins, has recently gained popularity as a launching platform for these tokens (Boaz Sobrado, 2025). It simplifies meme coin creation for non-technical users and features fair launch mechanisms through an initial bonding curve that raises token prices as purchases increase (pump.fun, 2025). In particular, a key characteristic of pump.fun tokens is their deterministic address format: all token addresses end with the string pump. This consistent naming convention provided a reliable method for identification. By leveraging this unique identifier, we discovered 4,140 meme coins, detecting a total of 28,975 different meme coins.
@@
By combining the known meme coins collected in Sec. 3.1.1 with the newly identified tokens, we obtain a total of 35,220 meme coins.
3.1.3. Dataset Refinement

### v2

@@ Abstract
From viral jokes to a billion-dollar phenomenon, meme coins have become one of the most popular segments in cryptocurrency markets. Unlike utility-focused crypto assets like Bitcoin, meme coins derive value primarily from community sentiment, making them vulnerable to manipulation. This study presents an unprecedented cross-chain analysis of the meme coin ecosystem, examining 34,988 tokens across Ethereum, BNB Smart Chain, Solana, and Base. We characterize their tokenomics and track their growth in a three-month longitudinal analysis. We discover that among high-return tokens (>100%), an alarming 82.8% show evidence of artificial growth strategies designed to create a misleading appearance of market interest. These include wash trading and a new form of manipulation we define as Liquidity Pool-Based Price Inflation (LPI), where small strategic purchases trigger dramatic price increases. We find that profit extraction schemes, such as pump and dumps and rug pulls, typically follow initial manipulations like wash trading or LPI, indicating how early manipulations create the foundation for later exploitation. We quantify the economic impact of these schemes, identifying over 17,000 victimized addresses with realized losses exceeding \$9.3 million. These findings reveal that combined manipulations are widespread among high-performing meme coins, suggesting that their dramatic gains are often driven by coordinated efforts rather than natural market dynamics.
1 Introduction
@@
The explosive rise of meme coins represents one of the most impressive trends in cryptocurrency markets. What started as a joke with Dogecoin in 2013 [nani2022doge] has evolved into a market worth over \$117 billion by 2024 [meme_success_cmc]. Their appeal lies in the potential for extreme returns, exemplified by Pepe (400,000%) and Shiba Inu (80,000,000%) [impressive_return_meme_coins_spa]. These extraordinary profits have attracted waves of investors hoping to find the next meme coin before its price explodes, leading to an unprecedented proliferation of meme coins, largely driven by platforms like pump.fun, which allow users to create tokens with minimal technical skills [pump_success]. The success of meme coins has prompted regulatory institutions to develop specialized definitions, underscoring their importance as a distinct asset. The U.S. Securities and Exchange Commission (SEC) defines meme coins as "a type of crypto asset inspired by internet memes, characters, current events, or trends for which the promoter seeks to attract an enthusiastic online community to purchase the meme coin and engage in its trading" [meme_coins_sec], a definition we adopt in this study. However, meme coins present a concerning vulnerability: Unlike utility tokens, meme coins rely exclusively on perceived interest to maintain their value [nani2022doge], creating strong incentives for manipulations.
@@
This study investigates whether token creators and early investors employ artificial growth strategies to create illusory activity to attract investors, and how these strategies facilitate subsequent profit extraction schemes like pump and dumps [xu2019anatomy] or rug pulls [zhou2024stop]. First, we collect and publicly release [meme_coin_repository] a dataset of 34,988 meme coins across four major blockchains (Ethereum, BNB Smart Chain (BSC), Solana, and Base), which, to our knowledge, is the largest curated dataset of meme coins to date. Our investigation begins with the characterization of meme coin tokenomics across different blockchains. We then perform a three-month longitudinal analysis of price growth, identifying 707 high-return tokens (>100%). We uncover an alarming 82.8% of these high-performing meme coins exhibit evidence of market manipulation, including recurrent wash trading and what we define as Liquidity Pool-Based Price Inflation (LPI), an operation where strategic purchases with minimal capital trigger dramatic price increases. Furthermore, we find profit extraction schemes, such as pump and dumps and rug pulls, designed to capitalize on artificially generated interest: 62.9% of tokens subjected to these operations had previously undergone wash trading or LPI, revealing how initial manipulations often serve as preliminary stages in broader exploitation strategies. This connection between manipulations is even stronger among meme coins delisted by established aggregators: 86.67% of the delisted tokens subjected to profit extraction had previously been artificially inflated. In such cases, token creators first use artificial growth tactics to meet listing requirements, then leverage their newfound visibility to execute exit strategies, maximizing their profits. We link these schemes to quantifiable financial harm, documenting \$3.27 million in realized losses from pump and dump schemes and \$6.04 million from rug pulls, affecting over 17,000 victim addresses. These findings highlight the prevalence of manipulation strategies among high-return meme coins, underscoring the need for enhanced investor protection mechanisms.
@@
  • Curated Dataset of Meme Coins: We build and publicly release [meme_coin_repository] the first large-scale dataset of 34,988 meme coins across Ethereum, BNB Smart Chain, Solana, and Base, providing a valuable resource for future works on the meme coins ecosystem.
  • Cross-chain Characterization: Leveraging this dataset, we conduct the first cross-chain analysis of meme coins, uncovering blockchain-specific patterns in tokenomics, temporal evolution, and price growth.
@@
  • Artificial Growth Detection: We identify strategies that create misleading market interest, including a new manipulation we define as Liquidity Pool-Based Price Inflation (LPI). We show that 82.8% of high-performing tokens exhibit manipulation indicators (e.g., concentrated ownership, wash trading, or LPI), suggesting that their growth is often due to coordinated manipulation.
  • Profit Extraction and Economic Impact: We show that profit extraction is often sequential, with 62.9% of high-return tokens (86.67% of delisted ones) using artificial growth to facilitate exit scams. We quantify the direct cost of these schemes, documenting over \$9.3 million in losses impacting over 17,000 addresses.
2 Related Work
@@
Memes and Meme Coins. Internet memes, defined by Shifman [shifman2013memes] as “units of popular culture that are circulated, imitated, and transformed by individual Internet users,” have become significant cultural phenomena. The intersection of meme culture and cryptocurrencies created meme coins, which differ from traditional cryptocurrencies through their community-driven nature, humor incorporation, and meme culture origins [stencel2023meme, nani2022doge]. Kim et al. [kim2024identifying] studied how social media drives meme coin valuation, while Belcastro et al. [belcastro2023enhancing] highlighted that meme coin price movements require specialized prediction approaches, different from other crypto assets. Long et al. introduced CoinCLIP [long2024coinclip], a framework leveraging visual-language models to classify meme coin viability, and released the Coin-Meme dataset [long2024bridging] containing multimodal data from Solana meme coins.
@@
Market Manipulations on Blockchains. The cryptocurrency ecosystem, characterized by its relative lack of regulatory oversight, has become susceptible to various forms of market manipulation [gandal2018price, krafft2018experimental, cernera2023ready, daian2020flash, cernera2024blockchain]. These practices undermine market integrity and pose significant risks to investors.
@@
Wash Trading is a manipulation that aims to create an illusion of market activity by generating artificial volume. Victor and Weintraud [victor2021detecting] found that up to 35% of ERC-20 token volume on some exchanges was generated through wash trading, while Le Pennec et al. [le2021wash] proposed machine learning approaches to identify wash trading activities involving ERC-20 tokens. This manipulation has also extended to non-fungible token (NFT) markets, with [von2022nft, la2023game] documenting significant wash trading activity in NFT marketplaces.
@@
Pump and dumps represent one of the most prevalent manipulations in cryptocurrency markets. These schemes involve inflating an asset’s price through misleading statements and coordinated buying (the "pump"), followed by selling the overvalued assets (the "dump") once enough investors have been lured in [kamps2018moon]. Xu and Livshits [xu2019anatomy] developed detection mechanisms for pump and dumps, while other studies [hamrick2019economics, la2023doge, la2020pump] found that pump announcements typically generated average returns of 65% for organizers, mostly at the expense of unsuspecting investors. These works also examined the communication platforms used to coordinate such schemes, identifying Telegram and Discord as the main hubs, a finding also supported by Telegram-focused research [la2023sa, la2018pretending, la2025tgdataset].
@@
Lastly, rug pulls have become a harmful exit scam in decentralized finance (DeFi), where developers abandon projects after attracting investors. Zhou et al. [zhou2024stop] analyzed 201 rug pulls and developed tools to detect associated risks, while Cernera et al. [cernera2023ready] identified "1-day rug pulls" that generated \$240 million in profits, primarily on BNB Smart Chain.
This work addresses several critical gaps in existing literature. First, we focus on meme coins, which possess unique characteristics that distinguish them from other cryptocurrency assets. These differences have led regulatory institutions like the SEC to develop specialized definitions for meme coins [meme_coins_sec], highlighting the need for dedicated analysis. Second, while previous studies examine individual manipulation types such as wash trading [la2023game, le2021wash] or pump-and-dumps [xu2019anatomy, la2020pump], we analyze how these different strategies are combined and executed sequentially to maximize profits. Finally, we provide a comprehensive cross-chain analysis of the meme coin ecosystem across four major blockchains (Ethereum, BSC, Solana, and Base), surpassing the scope of previous works, which examine at most two blockchains, and offering a more complete overview of market manipulations.
3 Methodology
@@
To capture a comprehensive view of the meme coin ecosystem, we design a methodology comprising three steps. We first identify tokens labeled as meme coins by established platforms, then expand our dataset by collecting newly created tokens from additional sources, and finally apply classification techniques to distinguish meme coins from other assets.
3.1 Data Collection
@@
The meme coin ecosystem spans multiple blockchains, including well-known tokens and emerging projects. To capture this diversity, we adopt a blockchain-agnostic data collection strategy, sourcing information from mainstream crypto tracking platforms and specialized decentralized exchange (DEX) aggregators. This cross-chain approach enables us to track migration patterns and observe how meme coin creators leverage different blockchains, uncovering cross-chain trends and insights that would be missed in a single-chain analysis.
3.1.1 Known Meme Coins
@@
We start with CoinMarketCap [coinmakretcap] and CoinGecko [coingecko], two leading cryptocurrency tracking platforms that maintain specialized meme token categories. CoinGecko defines meme coins as “Tokens with no intrinsic value, but with a strong community that fosters social interaction” [meme_coins_definition_coingecko], while CoinMarketCap describes them as “a type of cryptocurrency that originated from internet memes. They are often created as a joke or for the purpose of satire… Some of these coins have gained popularity due to their community-driven nature and the humorous themes they embody” [meme_coins_definition_coinmarketcap]. Both definitions align with the one provided by the SEC [meme_coins_sec], which we adopt as the formal definition for our work. Using their free APIs [coinmakretcap_api, coingecko_api], we collect data on 2,930 and 2,107 meme coins, respectively. Additionally, we utilize Gecko Terminal [geckoterminal], which categorizes tokens by meme trends. This approach reflects the definition of meme coins as assets inspired by Internet trends and culture [meme_coins_sec]. We develop a custom scraper to collect data on 4,735 meme coins from pools tagged with relevant meme categories, including animal, dog, cat, inu, and pump.fun (a platform for meme coin creation). By aggregating the results of these services and removing duplicates by identifying each token using the tuple (token_contract_address,blockchain)(token\_contract\_address,blockchain), we obtain a total of 8,852 distinct meme coins.
3.1.2 Searching other Meme Coins
@@
To expand our dataset, we implement a second approach focusing on newly created tokens. We develop scrapers for two DEX aggregators: DexScreener [dexscreener] and CoinSniper [coinsniper], which track real-time data on emerging tokens across decentralized exchanges. These platforms differ from CoinMarketCap and CoinGecko in their listing criteria. While the latter require stricter conditions, such as functional websites and listing on integrated exchanges [coingecko_listing, coinmarketcap_listing], DEX aggregators implement more permissive policies: DexScreener automatically indexes tokens immediately upon the creation of the liquidity pool and the initial transaction activity [dexscreener_listing], while CoinSniper requires only a basic application process, after which the projects appear immediately on their New Listings Page without further verification [coinsniper_listing]. Our scrapers collected 30,771 tokens from DexScreener and 44,611 from CoinSniper, resulting in 65,021 unique tokens.
@@
Although these platforms are commonly used to monitor meme coins due to their focus on newly launched tokens on DEXs, they could list a wide range of token types. To distinguish meme coins from other digital assets, we strictly adhere to the formal definition established by the SEC, which characterizes them as "crypto assets inspired by internet memes, characters, current events, or trends" [meme_coins_sec]. Our analysis revealed that these cultural and thematic associations are consistently embedded within token names, creating distinctive linguistic patterns that can be systematically identified. Using this insight, we developed a name-based classification approach to detect meme coins within our larger dataset.
@@
Name-Based Classification. We begin by analyzing the names of verified meme coins collected in Sec. 3.1.1 to identify common keywords that characterize meme tokens. To extract these keywords, we apply Term Frequency–Inverse Document Frequency (TF-IDF) analysis to our verified meme token dataset. This technique identifies distinctive terms by assigning higher weights to words that appear frequently in meme coin names but are uncommon in general cryptocurrency naming [ramos2003using]. Our methodology proceeds as follows: First, we preprocess token names by converting them to lowercase and removing special characters. Next, we tokenize the names into individual words (unigrams) and compute their TF-IDF scores. To determine the optimal number of keywords to retain, we apply the elbow method, which identifies the point where additional keywords provide diminishing marginal value [bholowalia2014ebk]. As shown in Fig. 7 in the Appendix, the elbow analysis suggests 150 keywords as the optimal cutoff point. We then manually review these candidates to remove general cryptocurrency terms (such as SOL or ETH) that appear due to cross-chain versions of the same token (e.g., , Dogecoin (SOL)). This refinement yields 126 meme-specific keywords. The resulting keyword set reflects the market’s thematic dominance: Animal-related terms (dog: 7.24%, cat: 6.92%, inu: 4.82%) and trend-based keywords (AI: 4.96%) are heavily represented. We apply this name-based classification to the token dataset collected in Sec. 3.1.2, identifying tokens as meme coins if their names contain one or more of the 126 meme-specific keywords. This approach detects 26,557 meme coins.
Pump.fun-Based Classification. In addition to keyword matching, we directly include tokens originating from the pump.fun platform (identifiable by the .pump suffix in their addresses). Two factors justify this platform-specific inclusion. First, pump.fun explicitly markets itself as a specialized platform for meme coin creation [pump_fun_platform], establishing a high baseline probability that its listed assets are meme-related. Second, to empirically validate this assumption, we checked manually 100 randomly sampled pump’ tokens. The inspection confirmed that 100% of the sampled tokens featured names referencing internet memes, pop culture, or viral trends, thereby strictly aligning with the SEC’s definition [meme_coins_sec]. In particular, this manual verification revealed also niche memes not captured by the name-based classification. In this way, we mitigate false negatives and ensure the inclusion of emerging trends. With this approach, we identified 4,140 meme coins, detecting overall 28,975 meme coins.
@@
By combining the known meme coins collected in Sec. 3.1.1 with the new tokens, we obtain 35,220 meme coins.
3.1.3 Dataset Refinement